You have looked at the current financial statements for Reigle Homes, Co. The company has an EBIT of $3,170,000 this year. Depreciation, the increase in net working capital, and capital spending were $241,000, $106,000, and $495,000, respectively. You expect that over the next five years, EBIT will grow at 15 percent per year, depreciation and capital spending will grow at 20 per year, and NWC will grow at 10 per year. The company currently has $18,300,000 in debt and 380,000 shares outstanding. After Year 5, the adjusted cash flow from assets is expected to grow at 3 percent indefinitely. The company’s WACC is 8.9 percent and the tax rate is 35 percent.

What is the price per share of the company's stock? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Share price $
Answer #1

Answer:

Share price = $78.15

Working:

1 2 $3,170,000 $241,000 15% $3,645,500 $2,369,575 20% $289,200 20% $594,000 10% $116,600 $1,948,175 15% $4,192,325 $2,725,011

The above excel with 'show formula' is as follows:

3 4 0.15 =D3*(1+E2) =E3*(1-35%) 0.2 =D6*(1+E5) 0.2 =D8*(1+E7) 0.1 =D10*(1+E9) =E4+E6-E8-E10 0.15 =E3*(1+F2) =F3*(1-35%) 0.2 =